OpenStudy (anonymous):

help help help help help help a bulldog does 72,000j of work in 48 seconds how much power does the bulldogs use?(show your work )

OpenStudy (jamesj):

I just answered this for you below

OpenStudy (anonymous):

below?

OpenStudy (jamesj):

The first time you asked this question. I'll copy my response.

OpenStudy (jamesj):

Power is Work/Time. Now we need to check units. The standard units of power is Watts = J/sec. So we're good shape. Just divide: Power = ( 72000 J ) / (48 seconds) = .... W

OpenStudy (anonymous):

i cant see it all it ends here. 48 seconds) = .... W

OpenStudy (jamesj):

Yes, I didn't write the division. You can do that.

OpenStudy (jamesj):

( 72000 J ) / (48 seconds) = (72000/48) W = ... W

OpenStudy (jamesj):

So what answer to you get?

OpenStudy (anonymous):

1500